OpenStudy (aye_my_peeps):

WHAT IS THE PERCENT DECREASE FROM 60 TO 39

OpenStudy (gottennis121):

Is this a multiple choice?

OpenStudy (aye_my_peeps):

NO :(

OpenStudy (janu16):

35%

OpenStudy (janu16):

The percent decrease we are going to use is "x" If something decreases by 30%, then you are actually multiplying the number by 70%, which is 100% - 30% = 70%. Therefore: 60 (1 - x) = 39 60 - 60x = 39 -60x = -21 x = 0.35 or 35% The percent decreased 35% from 60 to 39

OpenStudy (janu16):

hope it helps

OpenStudy (aye_my_peeps):

thnx @Janu16 and thnx@gottennis

OpenStudy (janu16):

no probs

rebeccaxhawaii (rebeccaxhawaii):

got it?